# Gauge-independent gravitational waves from a minimal dark $U(1)$ sector with viable dark matter candidates

### v1

@@ Abstract
Searches for stochastic gravitational wave backgrounds from first-order phase transitions offer a powerful probe of hidden sectors, but quantitative predictions in gauge theories are obstructed by the gauge dependence of the finite-temperature effective potential and the associated tunneling action. We study a minimal gauged $U(1)$ dark sector containing a dark Higgs and a dark photon, optionally supplemented by a vectorlike dark fermion, coupled to the Standard Model through Higgs portal or kinetic mixing. Using the Nielsen identity together with a controlled derivative expansion and power counting, we construct a gauge-independent effective action in the high- and low-temperature limits, enabling model-intrinsic nucleation dynamics and robust gravitational wave predictions. We perform dedicated Monte Carlo scans in both limits and map viable microscopic parameters to detector-facing peak frequencies and amplitudes, spanning bands relevant to pulsar timing arrays and planned space-based interferometers. In our scans, supercooled transitions typically produce the strongest signals, whereas parametrically high-temperature transitions are comparatively rare and tend to be weak. We further connect the phase transition phenomenology to viable dark matter candidates within the same minimal field content, providing benchmark targets for dark photon dark matter and dark fermion dark matter, and highlighting their complementarity with gravitational wave observables. Overall, our results provide an end-to-end, gauge-independent pipeline from a minimal hidden sector Lagrangian to gravitational wave spectra and cosmologically viable dark matter benchmarks, yielding the most reliable and concrete predictions to date for a minimal gauged $U(1)$ dark sector.
@@ 1 Introduction
The discovery of gravitational waves provides a novel probe of hidden sectors in which dark matter may reside. Searches for stochastic gravitational wave backgrounds produced by first-order phase transitions, either from the Standard Model (SM) augmented by new physics [1, 2, 3, 4, 5, 6, 7], or in a dark sector [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 33, 31, 32, 34, 35, 36], offer a powerful and complementary pathway to uncovering new physics, especially in light of the continuing non-detection of dark matter in conventional experiments. Gravitational wave observations thus constitute a key component of a multi-messenger program to test particle physics in the early Universe. Accordingly, the interplay between gravitational wave phenomenology and physics beyond the SM, particularly in scenarios aimed at explaining the nature of dark matter, has become one of the most active frontiers of current research.
In well-motivated dark sector models, gravitational wave from a first-order phase transition is particularly compelling: New gauge symmetries and scalar fields are ubiquitous in beyond the SM frameworks, and their symmetry breaking can naturally occur at scales both below and above the electroweak scale, leading to gravitational wave signals that span a wide range of frequencies. Consequently, it is essential to establish robust and concrete predictions for gravitational wave signals from dark sectors that can be confronted with data across multiple frequency bands. If detected, such a signal can reveal information about the underlying dark sector parameters including the gauge coupling, the scalar quartic coupling, and the scalar vacuum expectation value (vev), that may be difficult or impossible to access directly through collider probes or conventional dark matter searches.
@@
We focus in particular on supercooled first-order phase transitions, which are among the strongest and most distinctive sources of a stochastic gravitational wave background [48, 49, 50, 51, 52, 53, 54, 19, 20, 55, 56, 57, 58, 59, 21, 60, 61, 62, 63, 33, 34, 35, 64]. In such transitions, the nucleation and percolation temperatures satisfy $T_{n,p}\ll T_{c}$, which typically enhances the gravitational wave amplitude and shifts the peak frequency to lower values, often bringing the signal into the nanohertz band and thereby attracting substantial recent attention. Moreover, models that undergo strong supercooling often arise in nearly flat potentials, for which the tunneling action remains large and nucleation is delayed. In many such scenarios, the portal couplings to the SM are also small, making these models challenging to probe with collider searches or conventional dark matter experiments. Gravitational wave observations therefore become a leading discovery channel and can provide compelling evidence for this region of parameter space.
In this work we focus on the simplest representative setup: a minimal $U(1)_{x}$ dark sector featuring a dark Higgs field and a dark photon, and optionally a dark fermion, together with portal interactions connecting the hidden sector to the SM. This minimality is not merely aesthetic. It makes the parameter dependence transparent, allows for a systematic exploration of distinct thermal regimes, and admits dark matter candidates whose cosmological evolution can be computed consistently alongside the phase transition dynamics. In particular, we highlight two viable dark matter candidates already present in the minimal field content: (1) dark photon dark matter, which requires vanishing or negligibly small kinetic mixing; and (2) dark fermion dark matter, which relies on a non-negligible dark photon portal.
@@
The remainder of this paper is organized as follows. In Section 2 we introduce the minimal dark $U(1)_{x}$ sector and portal interactions that connect it to SM states. In Section 3 we review the gauge dependence of the finite-temperature effective potential and the construction of gauge-independent effective actions based on the Nielsen identity, with particular emphasis on the low-temperature regime relevant for supercooled phase transitions. In Section 4 we present the computation of gravitational wave signatures from first-order phase transitions, including the tunneling action and transition rate, characteristic temperatures, hydrodynamic parameters, and the resulting gravitational wave power spectra, and we compare gauge-dependent result and gauge-independent predictions. In Section 5 we analyze the cosmological evolution of dark sector particles, derive the coupled Boltzmann equations for the dark matter scenarios of interest, and present benchmark models and phenomenological implications. Finally, we summarize our conclusions and present an outlook in Section 6.
@@ 2.1 The $U(1)_{x}$ extension of the SM
where $\mathcal{L}_{{\rm SM}}$ and $\mathcal{L}_{{\rm hid}}$ is given by
@@
where $F_{\mu\nu},F_{x\,\mu\nu}$ are the field strengths of the hypercharge and $U(1)_{x}$ gauge fields respectively, $H$ is the SM Higgs doublet and $\Phi$ is the $U(1)_{x}$ Higgs field with $U(1)_{x}$ charge +1+1. $A_{\mu}^{\prime}$ is the $U(1)_{x}$ gauge field, and $\chi$ is a dark fermion with vector mass $m_{\chi}$ and $U(1)_{x}$ charge $Q_{x}=+1$. The most important parameters relevant for the study of both $U(1)_{x}$ gravitational wave physics and dark matter physics are: the $U(1)_{x}$ gauge coupling $g_{x}$, the $U(1)_{x}$ Higgs vev $v_{x}$ and its quartic coupling $\lambda_{x}$.
@@
$\mathcal{L}_{{\rm mix}}=\;-\frac{\delta}{2}F_{\mu\nu}F_{x}^{\mu\nu}-\lambda_{{\rm mix}}(H^{\dagger}H)(\Phi^{\ast}\Phi)^{2}\,$, (2.4)
where $\delta$ and $\lambda_{{\rm mix}}$ are mixing parameters, both receiving stringent constraints from experiments. The experimental constraint of $\lambda_{{\rm mix}}$ will be discussed in Section 5.2. The kinetic mixing can result in from either integrating out bifundamental fields charged under both the $U(1)_{Y}$ and $U(1)_{x}$, which can be absent; or from graviton loop effects, which are in general suppressed. We focus on two classes of dark matter candidates in this paper:
@@
$\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}$ $=\frac{1}{2}(\partial_{\mu}h_{x}+g_{x}A_{\mu}^{\prime}G)^{2}+\frac{1}{2}\bigl[\partial_{\mu}G-g_{x}A_{\mu}^{\prime}(\phi_{x}+h_{x})\bigr]^{2}$
@@
Together with the gauge fixing term and the ghost terms,
@@
where the masses of the longitudinal (time-like) gauge component $A^{\prime}_{0}$, the ghost and the Goldstone boson depend on the gauge parameter $\xi$.
@@ 2.2 Higgs portal through scalar mixing
We now discuss the Higgs portal interactions from $H,\Phi$ quartic mixing term. The full scalar potential reads
@@
An orthogonal transformation is performed to diagonalize the mass matrix, which transformed the original eigenbasis $(h^{0}\;\;h_{x}^{0})^{T}$ into mass eigenbasis $(h\;\;h_{x})^{T}$ as
@@
with $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$. The $U(1)_{x}$ gauge boson $A^{\prime}$ acquires a mass $M_{A^{\prime}}=g_{x}v_{x}$, referred to as a dark photon $\gamma^{\prime}$. The couplings of $\gamma^{\prime}$ to two Higgs fields can be obtained from the kinetic term
@@
The interactions of $h,h_{x}$ to SM particles are calculated to be
@@ 2.3 $U(1)$ portal through kinetic mixing
which diagonalizes the kinetic and mass matrices simultaneously and transform the original eigenbasis $V^{T}=(C,B,A^{3})$ to the mass eigenbasis $E^{T}=(A^{\prime},A_{\gamma},Z)$ with $E=\mathcal{R}^{T}V$. The resulting mass matrix is
@@
where $eQ_{i}$ is the electric charges of the SM fermions, and
@@ 3.1 The gauge-dependent scalar effective potential
$V_{{\rm eff}}(\phi_{c},T,\xi)=V_{0}(\phi_{c})+V_{0}^{{\rm\text{1-loop}}}(\phi_{c},\xi)+V_{{\rm T}}^{{\rm{\rm\text{1-loop}}}}(\phi_{c},T,\xi)+V_{{\rm daisy}}(\phi_{c},T,\xi)\,$, (3.1)
where $V_{0}$ is the tree-level potential, $V_{0}^{{\rm{\rm\text{1-loop}}}}$ and $V_{{\rm T}}^{{\rm{\rm\text{1-loop}}}}$ represent the one-loop corrections at zero and finite temperatures respectively, and $V_{{\rm daisy}}$ corresponds to the daisy contribution arising from higher-loop resummation. The tree-level potential $V_{0}$ is written as
@@
$=$ $\;\frac{1}{64\pi^{2}}\biggl[m_{h_{x}}^{4}\Bigl(\log\frac{m_{h_{x}}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)+m_{G}^{4}\Bigl(\log\frac{m_{G}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)$
$+3m_{A^{\prime}}^{4}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)-m_{c}^{4}\Bigl(\log\frac{m_{c}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)\biggr]\,.$ (3.6)
In the last line of the above equation, the coefficient 1-1 from the $m_{c}^{4}$ term arises from the +1+1 contribution of the longitudinal (time-like) gauge scalar mode, c.f., Eq. (2.9), together with the 2-2 contribution from the ghost.
@@
$V_{{\rm T}}^{{\rm 1-loop}}(\phi_{c},T)=\sum_{i}\pm\,g_{i}\,\frac{T^{4}}{2\pi^{2}}\,J_{B/F}\biggl(\frac{m_{i}^{2}(\phi_{c})}{T^{2}}\biggr)\,$, (3.7)
where the thermal function $J$ is defined as
@@ 3.2 Gauge-independent effective action
As discussed above, the effective potential evaluated at a fixed background field value is in general gauge-dependent, and therefore cannot by itself be used to define an unambiguous gravitational wave signal. A gauge-independent treatment can be formulated using the Nielsen identity, which states that an infinitesimal change in the gauge parameter $\xi$ can be compensated by a corresponding field redefinition of the background configuration. In this sense, the $\xi$ dependence of $V_{\rm eff}$ at fixed $\phi$ is a coordinate effect in field space. The induced shift of the background field from the field redefinition, is of higher order in the $\hbar$ expansion and therefore does not appear at the leading order.

### v2

@@ Abstract
Searches for stochastic gravitational wave backgrounds generated by first-order phase transitions offer a powerful probe of hidden sectors, but quantitative predictions in gauge theories are obstructed by the gauge dependence of the finite-temperature effective potential and the associated tunneling action. We study a minimal gauged $U(1)$ dark sector containing a dark Higgs and a dark photon, optionally supplemented by a vectorlike dark fermion, coupled to the Standard Model through the Higgs portal or kinetic mixing. Using the Nielsen identity together with a controlled derivative expansion and power counting, we construct a gauge-independent effective action in the high- and low-temperature limits, enabling model-intrinsic nucleation dynamics and robust gravitational wave predictions. We perform dedicated Monte Carlo scans in both limits and map viable microscopic parameters to detector-facing peak frequencies and amplitudes, spanning bands relevant to pulsar timing arrays and planned space-based interferometers. In our scans, supercooled phase transitions typically produce much stronger signals and are more likely to fall within the sensitivity range of current and future gravitational wave detectors, whereas parametrically high-temperature phase transitions generally yield weaker signals and are of less phenomenological interest. We further connect the phase transition phenomenology to viable dark matter candidates within the same minimal field content, providing benchmark targets for dark photon dark matter and dark fermion dark matter, and highlighting their complementarity with gravitational wave observables. Overall, our results provide an end-to-end, gauge-independent pipeline from a minimal hidden sector Lagrangian to gravitational wave spectra and cosmologically viable dark matter benchmarks, yielding the most reliable and concrete predictions to date for a minimal gauged $U(1)$ dark sector.
@@ 1 Introduction
The discovery of gravitational waves provides a novel probe of hidden sectors in which dark matter may reside. Searches for stochastic gravitational wave backgrounds produced by first-order phase transitions, either from the Standard Model (SM) augmented by new physics [1, 2, 3, 4, 5, 6, 7], or in a dark sector [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40], offer a powerful and complementary pathway to uncovering new physics, especially in light of the continuing non-detection of dark matter in conventional experiments. Gravitational wave observations thus constitute a key component of a multi-messenger program to test particle physics in the early Universe. Accordingly, the interplay between gravitational wave phenomenology and physics beyond the SM, particularly in scenarios aimed at explaining the nature of dark matter, has become one of the most active frontiers of current research.
In well-motivated dark sector models, gravitational waves from first-order phase transitions are particularly compelling: New gauge symmetries and scalar fields are ubiquitous in BSM frameworks, and their symmetry breaking can naturally occur at scales both below and above the electroweak scale, leading to gravitational wave signals that span a wide range of frequencies. Consequently, it is essential to establish robust and concrete predictions for gravitational wave signals from dark sectors that can be confronted with data across multiple frequency bands. If detected, such a signal can reveal information about the underlying dark sector parameters including the gauge coupling, the scalar quartic coupling, and the scalar vacuum expectation value (vev), that may be difficult or impossible to access directly through collider probes or conventional dark matter searches.
@@
We focus in particular on supercooled first-order phase transitions, which are among the strongest and most distinctive possible sources of a stochastic gravitational wave background [52, 53, 54, 55, 56, 57, 58, 19, 20, 59, 60, 61, 21, 62, 63, 22, 64, 65, 66, 34, 67, 35, 36, 38, 37, 68]. In such transitions, bubble nucleation and percolation occur well below the critical temperature, $T_{n,p}\ll T_{c}$. The false vacuum can dominate the energy density of the Universe for an extended period before the transition completes, leading to a large release of vacuum energy and typically enhancing the resulting gravitational wave amplitude. At the same time, the lower transition temperature and larger characteristic length scale shift the peak frequency to lower values, in some cases into the nanohertz band, thereby attracting substantial recent attention. For scenarios with small portal couplings to the SM, which may be difficult to probe through collider searches or conventional dark matter experiments, gravitational waves from supercooled phase transitions can provide a leading discovery channel for this otherwise elusive region of parameter space.
In this work, we focus on the simplest representative setup: a minimal $U(1)_{x}$ dark sector featuring a dark Higgs field and a dark photon, and optionally a dark fermion, together with portal interactions connecting the hidden sector to the SM. This minimality is not merely aesthetic. It makes the parameter dependence transparent, allows for a systematic exploration of distinct thermal regimes, and admits dark matter candidates whose cosmological evolution can be computed consistently alongside the phase transition dynamics. In particular, we highlight two viable dark matter candidates already present in the minimal field content: (1) dark photon dark matter, which requires vanishing or negligibly small kinetic mixing; and (2) dark fermion dark matter, which relies on a non-negligible dark photon portal.
@@
The remainder of this paper is organized as follows. In Section 2 we introduce the minimal dark $U(1)_{x}$ sector and portal interactions that connect it to SM states. In Section 3 we review the gauge dependence of the finite-temperature effective potential and the construction of gauge-independent effective actions based on the Nielsen identity, with particular emphasis on the low-temperature regime relevant for supercooled phase transitions. In Section 4 we present the computation of gravitational wave signatures from first-order phase transitions, including the tunneling action and transition rate, characteristic temperatures, hydrodynamic parameters, and the resulting gravitational wave power spectra, and we compare gauge-dependent results and gauge-independent predictions. In Section 5 we analyze the cosmological evolution of dark sector particles, derive the coupled Boltzmann equations for the dark matter scenarios of interest, and present benchmark models and phenomenological implications. Finally, we summarize our conclusions and present an outlook in Section 6.
@@ 2.1 The $U(1)_{x}$ extension of the SM
where $\mathcal{L}_{{\rm SM}}$ and $\mathcal{L}_{{\rm hid}}$ are given by
@@
where $F_{\mu\nu},F_{x\,\mu\nu}$ are the field strengths of the hypercharge and $U(1)_{x}$ gauge fields respectively, $H$ is the SM Higgs doublet and $\Phi$ is the $U(1)_{x}$ Higgs field with $U(1)_{x}$ charge +1+1. $A_{\mu}^{\prime}$ is the $U(1)_{x}$ gauge field, and $\chi$ is a dark fermion with vector mass $m_{\chi}$ and $U(1)_{x}$ charge $Q_{x}=+1$. The most important parameters relevant for the study of both $U(1)_{x}$ gravitational wave physics and dark matter physics are the $U(1)_{x}$ gauge coupling $g_{x}$, the $U(1)_{x}$ Higgs vev $v_{x}$ and its quartic coupling $\lambda_{x}$.
@@
$\mathcal{L}_{{\rm mix}}=\;-\frac{\delta}{2}F_{\mu\nu}F_{x}^{\mu\nu}-\lambda_{{\rm mix}}(H^{\dagger}H)(\Phi^{\ast}\Phi)\,$, (2.4)
where $\delta$ and $\lambda_{{\rm mix}}$ are mixing parameters, both receiving stringent constraints from experiments. The experimental constraint of $\lambda_{{\rm mix}}$ will be discussed in Section 5.2. The kinetic mixing can arise from either integrating out bifundamental fields charged under both the $U(1)_{Y}$ and $U(1)_{x}$, which can be absent; or from graviton loop effects, which are in general suppressed. We focus on two classes of dark matter candidates in this paper:
@@
$\bigl|(\partial_{\mu}-ig_{x}A_{\mu}^{\prime})\Phi\bigr|^{2}$ $=\frac{1}{2}(\partial_{\mu}h_{x}+g_{x}A_{\mu}^{\prime}G)^{2}+\frac{1}{2}\bigl[\partial_{\mu}G-g_{x}A_{\mu}^{\prime}(\phi_{c}+h_{x})\bigr]^{2}$
@@
Together with the gauge-fixing term and the ghost terms,
@@
where the masses of the unphysical time-like gauge component $A^{\prime}_{0}$, the ghost and the Goldstone boson depend on the gauge parameter $\xi$.
@@ 2.2 Higgs portal through scalar mixing
We now discuss the Higgs portal interactions from the quartic mixing term between $H$ and $\Phi$. The full scalar potential reads
@@
An orthogonal transformation is performed to diagonalize the mass matrix, which transforms the original eigenbasis $(h^{0}\;\;h_{x}^{0})^{T}$ into mass eigenbasis $(h\;\;h_{x})^{T}$ as
@@
with $c_{\theta}\equiv\cos\theta$ and $s_{\theta}\equiv\sin\theta$. The $U(1)_{x}$ gauge boson $A^{\prime}$ acquires a mass $M_{A^{\prime}}=g_{x}v_{x}$, referred to as a dark photon $\gamma^{\prime}$. The couplings of $\gamma^{\prime}$ to two Higgs fields can be obtained from
@@
The couplings of $h$ and $h_{x}$ to SM particles are then given by
@@ 2.3 $U(1)$ portal through kinetic mixing
which diagonalizes the kinetic and mass matrices simultaneously and transforms the original eigenbasis $V^{T}=(C,B,A^{3})$ to the mass eigenbasis $E^{T}=(A^{\prime},A_{\gamma},Z)$ with $E=\mathcal{R}^{T}V$. The resulting mass matrix is
@@
where $eQ_{i}$ are the electric charges of the SM fermions, and
@@ 3.1 The gauge-dependent scalar effective potential
$V_{{\rm eff}}(\phi_{c},T,\xi)=V_{0}(\phi_{c})+V_{0}^{{\rm\text{1-loop}}}(\phi_{c},\xi)+V_{T}^{{\rm{\rm\text{1-loop}}}}(\phi_{c},T,\xi)+V_{{\rm daisy}}(\phi_{c},T,\xi)\,$, (3.1)
where $V_{0}$ is the tree-level potential, $V_{0}^{{\rm{\rm\text{1-loop}}}}$ and $V_{T}^{{\rm{\rm\text{1-loop}}}}$ represent the one-loop corrections at zero and finite temperatures respectively, and $V_{{\rm daisy}}$ corresponds to the daisy contribution arising from higher-loop resummation. The tree-level potential $V_{0}$ is written as
@@
$=$ $\;\frac{1}{64\pi^{2}}\biggl[m_{h_{x}}^{4}\Bigl(\log\frac{m_{h_{x}}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)+3m_{A^{\prime}}^{4}\Bigl(\log\frac{m_{A^{\prime}}^{2}}{\Lambda^{2}}-\frac{5}{6}\Bigr)+m_{A_{0}^{\prime}}^{4}\Bigl(\log\frac{m_{A_{0}^{\prime}}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)$
$\qquad\,+m_{G}^{4}\Bigl(\log\frac{m_{G}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)-2m_{c}^{4}\Bigl(\log\frac{m_{c}^{2}}{\Lambda^{2}}-\frac{3}{2}\Bigr)\biggr]\,.$ (3.6)
where the contributions from the dark Higgs, the physical dark photon $A^{\prime}$ with three degrees of freedom, the unphysical scalar gauge mode $A_{0}^{\prime}$ with one degree of freedom, c.f., Eq. (2.9), the Goldstone boson, and the ghost are all included. The $m_{A_{0}^{\prime}}^{4}$ term combines with the $m_{c}^{4}$ term, and eventually the overall coefficient of the $m_{c}^{4}$ contribution becomes 1-1.
@@
$V_{T}^{{\rm 1-loop}}(\phi_{c},T)=\sum_{i}\pm\,g_{i}\,\frac{T^{4}}{2\pi^{2}}\,J_{B/F}\biggl(\frac{m_{i}^{2}(\phi_{c})}{T^{2}}\biggr)\,$, (3.7)
where the upper (lower) sign corresponds to bosons (fermions), and the thermal function $J$ is defined as
@@
In the conventional gravitational wave analysis, one starts from the full gauge-dependent effective potential in Eq. (3.1), and solves the bounce equation Eq. (4.5) in Landau gauge $\xi=0$. This leads to a gauge-dependent bounce solution, and consequently, gauge-dependent gravitational wave results, preventing definitive predictions for the minimal $U(1)$ dark sector. The results are nevertheless presented in Section 4.5.1, to enable a direct comparison with the gauge-independent analysis.
In the high-temperature regime, using the expansions in Eqs. (3.9) and (3.10), the finite-temperature effective potential can be conveniently approximated by the simple form
$V(\phi,T)=D\,(T^{2}-T_{0}^{2})\,\phi^{2}-E\,T\,\phi^{3}+\frac{\lambda_{x}}{4}\,\phi^{4}\,$, (3.16)
where $T_{0}$ is the temperature at which $\phi=0$ ceases to be a local minimum. A caveat is that the parametrization in Eq. (3.16) is valid only within the high-temperature regime and should not be applied at intermediate or low temperatures, as will be discussed in Section 4.5.1.
@@ 3.2 Gauge-independent effective action
As discussed above, the effective potential evaluated at a fixed background field value is in general gauge-dependent, and therefore cannot by itself be used to define an unambiguous gravitational wave signal. A gauge-independent treatment can be formulated using the Nielsen identity, which states that an infinitesimal change in the gauge parameter $\xi$ can be compensated by a corresponding field redefinition of the background configuration. In this sense, the $\xi$ dependence of $V_{\rm eff}$ at fixed $\phi$ is a coordinate effect in field space. The induced shift of the background field from the field redefinition is of higher order in the $\hbar$ expansion and therefore does not appear at the leading order.